OpenStudy (anonymous):

SLOPE 4x+y=7 help me get this into y=mx+b form, please?

geerky42 (geerky42):

Simply subtract both sides by 4x.

OpenStudy (anonymous):

4x + y = 7 subtract 4x from both sides 4x - 4x + y = 7 - 4x 0 + y = 7 - 4x y = 7 - 4x re write the order of the terms y = -4x + 7

geerky42 (geerky42):

^

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

Thanks, it's been years since simple algebra and my brain wanted to overthink it! lol I feel dumb.